# NEST: Self-supervised Fast Conformer as All-purpose Seasoning to Speech Processing Tasks

### v3

@@ Abstract
Self-supervised learning has been proved to benefit a wide range of speech processing tasks, such as speech recognition/translation, speaker verification and diarization, etc. However, most of current approaches are computationally expensive due to lacking in sub-sampling or using clustering based speech quantization. In this paper, we propose a simplified and more efficient self-supervised learning framework termed as NeMo Encoder for Speech Tasks (NEST). Specifically, we adopt the FastConformer architecture, which has an 8x sub-sampling rate and is faster than Transformer or Conformer architectures. Instead of clustering-based token generation, we resort to fixed random projection for its simplicity and effectiveness. We also propose a generalized noisy speech augmentation that teaches the model to disentangle the main speaker from noise or other speakers. Experiments show that NEST improves over existing self-supervised models on a variety of speech processing tasks. Code and checkpoints will be available via NVIDIA NeMo toolkit [^1] https://github.com/NVIDIA/NeMo.
@@ I Introduction
Most recent speech self-supervised models are inspired by the BERT [1] model, which learn text token embedding by predicting the target of the masked positions given the context of the unmasked ones. Among them, are two main streams of contrastive and predictive models. The contrastive approach [2, 3, 4, 5] quantizes the speech features into a set of target feature vectors and trains with a contrastive loss using the positive and negative target features. Meanwhile, the predictive approach [6, 7, 8, 9] quantizes the speech features into tokens and train with masked token prediction loss as in BERT [1]. In addition to the two approaches, some works also learn from the masked auto-encoding [10] approach and train speech self-supervised models with a reconstruction objective [11, 5].
@@
HuBERT [6], as a pioneer work of the predictive approach, generates the target tokens by running k-means clustering on the middle layer features extracted from anohter SSL model that is pretrained for a small number of steps. Then, W2v-BERT [14] proposes to combine the training objectives of both Wav2vec-2.0 [2] and HuBERT [6] by applying contrastive loss on the middle layer output while predictive loss at the final output layer. Later, BEST-RQ [7] shows that the clustering based token generation can be replaced by simple fixed random-projection quantization, and this simple modification is able to match or outperform HuBERT on ASR.
@@
In order to improve performance on speaker tasks, WavLM [8] proposes a noisy speech augmentation technique and a denoising masked token prediction objective, by adding a speech segment of a different speaker to the current speech and training the model to predict the target tokens generated using original clean speech. XEUS [9] further extends WavLM [8] by adding a de-reverberation task and extending to multilingual data of 1M hours.
However, previous SSL models also have their different limitations. First, some models [2, 6, 8] use CNN+transformer architecture which has a short frame length of 20ms, which slows down the model’s inference speed. Second, the HuBERT-style quantization is very computationally expensive, which could take 20% of total training time according to XEUS [9]. Third, although BEST-RQ [7] uses Conformer [15] encoder with 40ms frame length and simple random quantization, it lacks the ability to explicitly tell one speaker from another, which limits its performance on speaker tasks (e.g., speaker diarization).
In this paper, we tackle all these challenges and bring the best practices from previous works, which constitute the proposed NeMo Encoder for Speech Tasks (NEST) framework. Our contributions are summarized as follows:
  • A new speech SSL model that achieves SOTA performance with a simplified and more efficient framework.
  • Experiments show that NEST can help achieve SOTA performance on a variety of downstream tasks (ASR, AST, SLU, SD, etc).
  • Different from previous SSL works that focus mostly on downstream tasks with limited data, we also show that NEST can benefit speech recognition and translation even when data is relatively larger.
@@ II-C Speech Quantization
We resort to BEST-RQ [7] as our speech quantization method. Specifically, we use a single randomly initialized and frozen codebook of 8192 vocabulary and 16 dimension features. A randomly initialized and frozen linear layer is applied to the input Mel-spectrogram features to project them into the same dimension as the codebook, then a nearest neighbor search is applied to obtain the target tokens. Since there is a 8x subsampling, we channel-concatenate the features for each 8 consecutive frames before feeding into the linear layer, such that the lengths for the target tokens and input features are equal.
@@ III-A Dataset and Settings
We train the NEST-L (108M) and NEST-XL (600M) models using 100K hours of both English speech data, including 60K hours from LibriLight [22], 24K hours from English subset of Voxpopuli [23], and about 20K hours sampled data from the combination of Fisher [24], Switchboard [25], WSJ [26], NSC [27], People’s Speech [28]. The audios for speech augmentation is randomly selected within each batch, while we use noise audios from MUSAN [29] and Freesound [30]. We train the models with global batch size of 2048 for 80K steps on 128 NVIDIA A100 GPUs, with Noam annealing [16] and peak learning rate of 0.004, weight decay of 1e-3, gradient clipping 1.0 and warm-up of 25K steps. We set the speech augmentation probability as 0.2, among which we set noise and speech augmentation probabilities as 0.1 and 0.9.
@@ III-B Results on SUPERB Multi-task Speech Processing
As presented in Table I, our NEST-L model is able to outperform WavLM-base++ [8] with similar size of parameters on all tasks, and also outperforms WavLM-large [8] that is 3x as large on speaker verification (SV), speaker diarization (SD) and phoneme recognition (PR). When compared with the XEUS [9] model that is trained on 10x data, we can see that our NEST-XL model is still able to achieve better performance on all speaker and content tasks, with especially large improvements on speaker verification, speaker diarization and phoneme recognition. Overall, we are able to achieve new state-of-the-art results on SID, SV, SD, PR and ASR tasks compared with WavLM [8] that has data size as well as XEUS [9] that is trained on much large data, demonstrating the effectiveness of NEST when applied on various downstream speech processing tasks.
TABLE III: Results on speech translation from English to German, French and Spanish. BLEU score is used as the metric, while punctuation and capitalization are included in metric calculation. “*” indicates second best performance.
@@
Canary-1B [21] 1B 86K 32.53 40.84 30.65 23.83 35.73 28.28 32.15* 22.66* 40.76* 29.50 33.07 33.23
NEST-XL-Transformer 1B 42K 30.87* 39.95* 30.01 22.82* 34.92* 27.99* 29.50 22.61 39.27 27.73* 32.51* 32.42*
TABLE IV: DER results on speaker diarization. “*” indicates second best.
Model DIHARD3 eval CALLHOME-part2
@@
EEND-EDA [34, 35] 15.55 7.83* 12.29 17.59
NeMo MSDD [36] 29.40 11.41 16.45 19.49*
@@
NEST-L-MLP 16.83 7.88 11.71* 23.46
RandFC-L-Sortformer 18.93 9.39 13.56 23.30
NEST-L-Sortformer [sortformer] 16.06* 6.40 11.02 20.10
@@ III-C Results on Multi-lingual Speech Recognition
Besides multi-task evaluation, we also study if an SSL model trained on single language can help other languages. To this end, we train an ASR model on four different languages: English (En), German (De), French (Fr), Spanish (Es). Specifically, we train an ASR model using NEST-XL as weight initialization and the hybrid-CTC-RNNT loss [37]. The training data comprises of 2.5K hours of German speech (MCV, MLS, Voxpopuli), 8.5K hours of English speech (MCV [38], MLS [39], Voxpopuli [23], SPGI [40], Europarl [41], LibriSpeech [42], NSC1 [27]), 1.4K hours of Spanish speech (MCV, MLS, Voxpopuli, Fisher [24]) and 1.9K hours of French speech (MCV, MLS, Voxpopuli). For baselines, we train another model using an English ASR model [43] as weight initialization, and also include some of the best ASR models like Whisper [20], SeamlessM4T [19] and Canary [21]. We run all models with the same beam size 5 with no language models on test sets of MCV-16.1 [38] and Voxpopuli [23].
@@ III-E Results on Speaker Diarization
To evaluate our model’s performance on speaker diarization, we train two variants of end-to-end diarization models: (1) simple two-layer multi-layer-perceptron (MLP) on top of FastConformer encoder and train with PIL; (2) more sophisticated Sortformer [sortformer] model with 18 layers of transformer on top of the encoder and train with a hybrid PIL and sorting loss. We also apply both NEST and random initialization to both methods for comparison. For training data, we use a combination of real data (Fisher [24], AMI Mix-Headset [47], ICSI [48], DIHARD3 dev [49], VoxConverse v0.3 [50], AISHELL [51]) and simulated data (composed from LibriSpeech [42], SRE [52, 53]) generated by the NeMo speech data simulator [54], with a total of 6.4K hours. We evaluate models’ performance on DIHARD3-eval [49] and CALLHOME-part2222We used splits from the Kaldi x-vector recipe [55]: Part 1 for training and Part 2 for evaluation [56].
As in Table IV, by comparing RandFC-L-MLP with NEST-L-MLP, and RandFC-L-Sortformer with NEST-L-Sortformer, we can see that NEST is able to significantly reduce the DER of models using randomly initialized encoder, showing the importance of using NEST in speaker diarization. We can also notice that Sortformer [sortformer] with NEST initialization is able to achieve second best results on DIHARD3 eval set, and outperforms EEND-EDA [34] on 2 and 3 speaker settings of the CALLHOME-part2.
TABLE V: Results on SLURP [57] benchmark for speech joint intent detection and slot filling.
@@ III-F Results on Spoken Language Understanding
As we can see, among all SSL-based SLU models, the proposed NEST model achieves the best performance on both intent detection accuracy and slot filling F1 scores. We also notice that scaling up from NEST-L to NEST-XL does bring some improvement on precision score on slot filling, but do not have significant effects on other metrics. In addition, compared with the NeMo-SSL-FC-Trans-L [62] baseline, we can see a more than 2% absolute improvement on F1 score by merely replacing the SSL speech encoder with NEST while keeping other hyper-parameters the same, which demonstrate the instant benefits that NEST can bring to existing speech processing models.
@@ IV Conclusion
In this paper, we propose a speech self-supervised framework termed NEST, and extensive experiments on multiple speech processing tasks show that the NEST model can help achieve state-of-the-art performance. Code, configurations and checkpoints are available through NVIDIA NeMo toolkit.

### v4

@@ Abstract
Self-supervised learning (SSL) has been proved to benefit a wide range of speech processing tasks, such as speech recognition/translation, speaker verification and diarization, etc. However, most of current speech SSL approaches are computationally expensive. In this paper, we introduce a simplified and more efficient SSL framework, termed as NeMo Encoder for Speech Tasks (NEST). Specifically, we adopt the FastConformer architecture with 8x sub-sampling rate, which is faster than Transformer or Conformer architectures. Instead of clustering-based quantization, we use fixed random projection for its simplicity and effectiveness. We also implement a generalized noisy speech augmentation that teaches the model to disentangle the main speaker from noise or other speakers. Experiments show that NEST improves over existing self-supervised models and achieves new state-of-the-art performance on a variety of speech processing tasks, such as speech recognition/translation, speaker diarization, spoken language understanding, etc. Code and checkpoints will be publicly available via NVIDIA NeMo framework [^1] https://github.com/NVIDIA/NeMo.
@@ I Introduction
Most recent speech self-supervised models are inspired by the BERT [1] model, which learn text token embedding by predicting the target of the masked positions given the context of the unmasked ones. Among them are two main streams of contrastive and predictive models. The contrastive approach [2, 3, 4, 5] quantizes the speech features into a set of target feature vectors and trains with a contrastive loss using the positive and negative target features. Meanwhile, the predictive approach [6, 7, 8, 9] quantizes the speech features into tokens and train with masked token prediction loss as in BERT [1]. In addition to the two approaches, some works also learn from the masked auto-encoding [10] approach and train speech self-supervised models with a reconstruction objective [11, 5].
@@
Figure 1: NEST serves as a bird nest that incubates the variety of speech task models.
HuBERT [6], as a pioneer work of the predictive approach, generates the target tokens by running k-means clustering on the middle layer features extracted from anohter SSL model that is pretrained for a small number of steps. Then, W2v-BERT [14] proposes to combine the training objectives of both Wav2vec-2.0 [2] and HuBERT [6] by applying contrastive loss on the middle layer output and predictive loss on the final output layer. Later, BEST-RQ [7] shows that the clustering based token generation can be replaced by simple fixed random-projection quantization, and this simple modification is able to match or outperform HuBERT on ASR.
@@
In order to improve performance on speaker tasks, WavLM [8] proposes a noisy speech augmentation technique and a denoising masked token prediction objective, by adding a speech segment of a different speaker to the current speech and training the model to predict the target tokens generated using original clean speech. XEUS [9] further extends WavLM [8] by adding a de-reverberation task and trains on multilingual data of 1M hours.
However, previous SSL models have notable limitations. First, several models [2, 6, 8] employ a CNN-Transformer architecture with a relatively short frame length of 20 ms, which negatively impacts inference speed. Second, HuBERT-style quantization is highly computationally intensive, consuming up to 20% of the total training time, as reported by XEUS [9]. Third, although BEST-RQ [7] uses Conformer [15] encoder with 40ms frame length and simple random quantization, it lacks the ability to explicitly tell one speaker from another, which limits its performance on speaker tasks such as speaker diarization.
In this paper, we tackle all these challenges and bring the best practices from previous works, which constitute the proposed NeMo Encoder for Speech Tasks (NEST). Our contributions are summarized as follows:
  • A new speech self-supervised learning framework with a simplified and more streamlined design.
  • Experiment results demonstrate that NEST can help achieve SOTA performance on a variety of downstream tasks (ASR, AST, SLU, SD, etc).
  • Unlike previous SSL approaches that primarily focus on downstream tasks with limited data, we also show that NEST can benefit speech recognition and translation even when data is relatively larger.
@@ II-C Speech Quantization
We use BEST-RQ [7] for speech quantization. Specifically, we employ a single randomly initialized and frozen codebook of 8192 vocabulary and 16 dimension features. A randomly initialized and frozen linear layer is applied to the input Mel-spectrogram features to project them into the same dimension as the codebook, then a nearest neighbor search is applied to obtain the target tokens. Since there is a 8x subsampling, we channel-concatenate the features for each 8 consecutive frames before feeding into the linear layer, such that the lengths for the target tokens and input features are equal.
@@ III-A Dataset and Settings
We train the NEST-L (108M) and NEST-XL (600M) models using 100K hours of English speech data, including 60K hours from LibriLight [22], 24K hours from English subset of Voxpopuli [23], and about 20K hours sampled data from the combination of Fisher [24], Switchboard [25], WSJ [26], NSC [27], People’s Speech [28]. The audios for speech augmentation are randomly selected within each batch, while we use noise audios from MUSAN [29] and Freesound [30]. We train the models with global batch size of 2048 for 80K steps on 128 NVIDIA A100 GPUs, with Noam annealing [16] and peak learning rate of 0.004, weight decay of 1e-3, gradient clipping 1.0 and warm-up of 25K steps. We set the speech augmentation probability as 0.2, among which we set noise and speech augmentation probabilities as 0.1 and 0.9.
@@ III-B Results on SUPERB Multi-task Speech Processing
As presented in Table I, our NEST-L model is able to outperform WavLM-base++ [8] with similar size of parameters on all tasks, and also outperforms WavLM-large [8] that is 3x as large on speaker verification (SV), speaker diarization (SD) and phoneme recognition (PR). When compared with the XEUS [9] model that is trained on 10x data, we can see that our NEST-XL model is still able to achieve better performance on all speaker and content tasks, with especially large improvements on speaker verification, speaker diarization and phoneme recognition. Overall, we are able to achieve new state-of-the-art results on SID, SV, SD, PR and ASR tasks compared with WavLM [8] that has similar data size as well as XEUS [9] that is trained on much large data, demonstrating the effectiveness of NEST when applied on various downstream speech processing tasks.
TABLE III: Results on speech translation from English to German, French and Spanish. BLEU score is used as the metric, while punctuation and capitalization are included in metric calculation. Underline indicates second best performance.
@@
Canary-1B [21] 1B 86K 32.53 40.84 30.65 23.83 35.73 28.28 32.15 22.66 40.76 29.50 33.07 33.23
[1.5ex] NEST-XL-Transformer 1B 42K 30.87 39.95 30.01 22.82 34.92 27.99 29.50 22.61 39.27 27.73 32.51 32.42
TABLE IV: DER results on speaker diarization. Underline indicates the second best evalutations. Starred(*) systems are not end-to-end systems which involve clustering steps.
Model DIHARD3-eval CALLHOME-part2
@@
EEND-EDA [34, 35] 15.55 7.83 12.29 17.59
WavLM-L+EEND-VC [8] N/A 6.46 10.69 11.84
EEND-GLA-small [35] 14.39 6.94 11.42 14.49
NeMo MSDD [36] 29.40 11.41 16.45 19.49
@@
NEST-L-MLP 16.83 7.88 11.71 23.46
RandFC-L-Sortformer-HL [37] 18.93 9.39 13.56 23.30
NEST-L-Sortformer-HL [37] 16.06 6.40 11.02 20.10
NEST-L-Sortformer-HL-PP [37] 14.60 6.08 9.57 15.40
@@ III-C Results on Multi-lingual Speech Recognition
Besides multi-task evaluation, we also study if an SSL model trained on single language can help other languages. To this end, we train an ASR model on four different languages: English (En), German (De), French (Fr), Spanish (Es). Specifically, we train an ASR model using NEST-XL as weight initialization and the hybrid-CTC-RNNT loss [38]. The training data comprises of 8.5K hours of English speech (MCV [39], MLS [40], Voxpopuli [23], SPGI [41], Europarl [42], LibriSpeech [43], NSC1 [27], Fisher [24]), 2.5K hours of German speech (MCV, MLS, Voxpopuli), 1.4K hours of Spanish speech (MCV, MLS, Voxpopuli) and 1.9K hours of French speech (MCV, MLS, Voxpopuli). For baselines, we train another model using an English ASR model [44] as weight initialization, and also include some of the best ASR models like Whisper [20], SeamlessM4T [19] and Canary [21]. We run all models with the same beam size 5 with no language models on test sets of MCV-16.1 [39] and Voxpopuli [23].
@@ III-E Results on Speaker Diarization
To assess the impact of NEST on speaker diarization, we train two variants of end-to-end diarization models: (1) simple two-layer multi-layer-perceptron (MLP) on top of FastConformer encoder and train with PIL; (2) more sophisticated Sortformer [37] hybrid loss (HL) model with post processing (PP) [^2] Post-processing parameters were tuned separately for DIHARD3 and CALLHOME on corresponding training parts. with 18 layers of transformer on top of the encoder. We also apply NEST and random initialization to both models for comparison. For training data, we use a combination of 2030 hours of real data (Fisher English [24], AMI Mix-Headset train+dev [48], ICSI [49], DIHARD3 dev [50], VoxConverse v0.3 [51], AISHELL-4 [52], CALLHOME-part1333We follow splits from the Kaldi x-vector recipe [53] by using part1 for training and part2 for evaluation. [54]) and 5150 hours of simulated data (composed from LibriSpeech [43] and SRE [55, 56]) generated by the NeMo speech data simulator [57]. We evaluate models’ performance on DIHARD3-eval [50] and CALLHOME-part2 [54].
As shown in Table IV, by comparing RandFC-L-MLP with NEST-L-MLP, and RandFC-L-Sortformer-HL with NEST-L-Sortformer-HL, we can see that NEST provides significant improvements ($1\sim 5\%$ absolute DER in different settings) over randomly initialized encoder, which demonstrates the effectiveness of NEST in speaker diarization task. We can also see that Sortformer [37] with NEST initialization is able to achieve second best results on DIHARD3 eval set when using postprocessing (PP), and it also achieves new SOTA results on 2 and 3 speaker settings of CALLHOME-part2 within all compared methods. Among end-to-end methods, NEST-L-Sortformer-HL-PP is able to outperform EEND-EDA [34] on all test sets, while RandFC-L-Sortformer-HL lags behind, showing that NEST is essential for achieving SOTA results in end-to-end speaker diarization.
TABLE V: Results on SLURP [58] benchmark for end-to-end speech joint intent detection and slot filling.
@@ III-F Results on Spoken Language Understanding
As we can see, among all SSL-based SLU models, using NEST as speech encoder can help achieve the best performance on both intent detection accuracy and slot filling F1 scores. We also notice that scaling up from NEST-L to NEST-XL does bring some improvement on precision score on slot filling, but do not have significant effects on other metrics. In addition, compared with the NeMo-SSL-FC-Trans-L [63] baseline, we can see a more than 2% absolute improvement on F1 score by merely replacing the SSL speech encoder with NEST while keeping other hyper-parameters the same, which demonstrates the instant benefits that NEST can bring to existing speech processing models.
@@ IV Conclusion
In this paper, we introduced a simplified and efficient speech self-supervised learning framework termed NEST, and extensive experiments on multiple speech processing tasks show that the NEST framework can help achieve state-of-the-art performance. Code, configurations and checkpoints are available through NVIDIA NeMo framework.